### v1

# Multispectral State-Space Feature Fusion: Bridging Shared and Cross-Parametric Interactions for Object Detection

Abstract
Modern multispectral feature fusion for object detection faces two critical limitations: (1) Excessive preference for local complementary features over cross-modal shared semantics adversely affects generalization performance; and (2) The trade-off between the receptive field size and computational complexity present critical bottlenecks for scalable feature modeling. Addressing these issues, a novel Multispectral State-Space Feature Fusion framework, dubbed MS2Fusion, is proposed based on the state space model (SSM), achieving efficient and effective fusion through a dual-path parametric interaction mechanism. More specifically, the first cross-parameter interaction branch inherits the advantage of cross-attention in mining complementary information with cross-modal hidden state decoding in SSM. The second shared-parameter branch explores cross-modal alignment with joint embedding to obtain cross-modal similar semantic features and structures through parameter sharing in SSM.
Finally, these two paths are jointly optimized with SSM for fusing multispectral features in a unified framework, allowing our MS2Fusion to enjoy both functional complementarity and shared semantic space. Benefiting from the design of the dual-branch SSM, our approach simultaneously inherits the computational efficiency and the global receptive field, significantly improving the performance of multispectral object detection. In our extensive experiments on mainstream benchmarks including FLIR, M3FD and LLVIP, our MS2Fusion significantly outperforms other state-of-the-art multispectral object detection methods, evidencing its superiority. Moreover, MS2Fusion is general and applicable to other multispectral perception tasks. We show that, even without specific design, MS2Fusion achieves state-of-the-art results on RGB-T semantic segmentation and RGB-T salient object detection, showing its generality. The source code will be available at https://github.com/61s61min/MS2Fusion.git.
keywords:
Multispectral Object Detection, State Space Model, Shared-Parameter, Cross-Parameter Interaction
journal: Nuclear Physics B
1 Introduction
Multispectral object detection has recently drawn increasing interest owing to its robust performance by fusing information from multiple spectral bands, such as RGB and thermal bands. RGB images usually offer high resolution, rich color and texture features, but suffer from sharply performance deterioration in complex scenarios such as low light, adverse weather or occlusions. In contrast, thermal images can effectively overcome these environmental limitations but exhibit obvious deficiencies in color and texture details. Current single-modal generic object detection methods struggle to overcome these aforementioned challenges. However, multispectral feature fusion paves a way to provide a reliable object detection solution under such challenging conditions.
As shown in Figure 1a, existing studies generally suggest that complementary features play a critical role when one modality is insufficient. For example, RGB images provide discriminative color and texture cues when thermal objects lack distinct contours, while thermal images offer thermal signatures when RGB imaging suffers from low illumination or occlusion. However, as demonstrated in Figure 1b, scenarios where both modalities exhibit weak discriminative features (e.g., blurred textures in RGB and low contrast in thermal) cannot be resolved by complementary information alone. In such cases, shared features, such as cross-modal consistent shapes and structural patterns, become essential, as they capture modality-invariant representations for reliable detection. Thus, we speculate that a robust multispectral object detection framework should dynamically leverage both complementary and shared features to handle diverse real-world challenges.
Figure 1: The pros and cons of RGB (left) and thermal (right) images. (a) Both modalities provide complementary information, and their fusion enables more robust object detection; (b) Dual-modal shared features become crucial, since neither modality stands out distinctly. (e.g., modality-specific characteristics such as texture and thermal radiation are blurred, and cross-modal consistent features like object contours and structures are helpful for detection.)
Previous studies [1, 2, 3, 4, 5, 6, 7] have predominantly focused on complementary feature learning across modalities, often overlooking the exploration of shared feature representations or inherent structural similarities between them. Moreover, existing approaches typically employ a single fusion strategy to directly combine multi-modal inputs, neglecting the potential benefits of adaptive or hierarchical fusion mechanisms. These methods do not fully explore or exploit cross-modal shared features, ignoring the potential effect of enhancing the performance of single-modal features. This fusion paradigm often suppresses weaker yet discriminative features during cross-modal integration, leading to significant information loss. For multispectral object detection, shared feature representation plays a pivotal role in multi-modal fusion. Not only does it mitigate cross-modal discrepancies, but it also augments single-modal features, thereby substantially improving feature expressiveness and detection robustness in challenging environments.
In addition, most of the mainstream methods leverage CNN [1] or Transformer [8, 9, 2] for feature fusion. Albeit effective, existing CNN-based methods often struggle with capturing broader contextual information across modalities due to their limited receptive fields. On the other hand, despite excellence at modeling global dependencies, the Transformer-based approaches may degrade with longer input sequences, leading to performance degradation and higher computational costs as model complexity. These factors restrict their practicality in resource-constrained environments, underscoring the need for more efficient fusion strategies in multispectral object detection.
Figure 2: Comparing Transformer-based fusion (a), Mamba-based fusion (b) and our proposed MS2Fusion (c), with $F_{T}$ and $F_{V}$ as the input thermal and RGB image features, respectively. In Transformer-based method (a), $F_{T}$ and $F_{V}$ are fused through the multi-head attention mechanism, effectively integrating complementary information and enhancing performance across scenarios. The traditional Mamba approach (b) directly mixes dual-modal features to generate B, C, and $\Delta$ parameters for SSM-based feature interaction, which may lead to modal misalignment and feature redundancy. In contrast, our method (c) first performs intra-modal feature interaction and then extracts cross-modal shared features, achieving better modal alignment and fusion, thereby providing more robust and unified feature representations.
(a) CNN-based fusion
(b) Transfomer-based fusion
(c) MS2Fusion
Figure 3: Effective receptive field visualizations comparing CNN-based fusion method (a), Transformer-based fusion method (b), and the proposed MS2Fusion (c) method. Quantitative analysis demonstrates that MS2Fusion achieves significantly broader receptive field coverage compared to the others.
Recent advances in sequence modeling show SSM-based methods excel by compressing features into compact hidden states, enabling efficient inference with constant-time full-sequence processing. Mamba [10] enhances this with selective state spaces, dynamically retaining task-relevant features. Vision Mamba (Vim) [11] further proves its effectiveness for visual tasks, boosting both efficiency and performance.
Inspired by this, we propose MS2Fusion, a novel framework that simultaneously leverages complementary and shared features across modalities while overcoming limitations of CNN and Transformer in multispectral feature fusion. Figure 2 reveals that Transformer-based methods (a) and existing Mamba solutions (b) fail to exploit cross-modal shared features. In contrast, our method (c) explicitly models both complementary and shared cross-modal interactions through three key components:
  • 1.
    Cross-Parametric State Space Model (CP-SSM): Facilitates implicit feature complementarity by exchanging output matrices between modality-specific state spaces, enabling cross-modal feature enrichment while preserving modality-specific characteristics.
  • 2.
    Shared-Parametric State Space Model (SP-SSM): Learns a unified feature space through parameter sharing, aligning heterogeneous modality distributions to extract discriminative shared representations that enhance single-modality features.
  • 3.
    Feature Fusion State Space Model (FF-SSM): Introduces a bidirectional input scheme to expand the Effective Receptive Field (ERF) of state spaces, mitigating feature attenuation while enabling adaptive fusion of cross-modal information.
As demonstrated in Figure 3, our ERF analysis reveals that MS2Fusion achieves superior spatial coverage compared to both CNN and Transformer, successfully integrating local details with global context. This capability addresses the fundamental constraint of standard CNN in modeling long-range dependencies while maintaining computational efficiency.
Table 1 also provides our component replacement experiments on the MS2Fusion module in the multispectral object detection framework (Figure 5), which yield three key findings: (1) Compared to Transformer, MS2Fusion achieves a 66.6% reduction in FLOPs and 70.4% fewer parameters while delivering better detection accuracy (+8.3 mAP50 points); (2) When benchmarked against CNN baseline, it maintains 26.0% lower computational costs and 18.4% parameter reduction, while achieving significant improvements of +9.0 mAP50 points; (3) The MS2Fusion consistently outperforms both baselines across all complexity-accuracy trade-off metrics. These systematic experiments provide conclusive evidence that MS2Fusion successfully breaks the traditional efficiency-accuracy trade-off, establishing new state-of-the-art performance in multispectral object detection.
Table 1: Comparison of different multispectral feature fusion methods
GFLOPs Params (M) mAP@0.5 mAP@0.75 mAP
CNN 190.3 159.7 74.3 23.8 32.5
Transformer 421.9 440.6 75.0 24.1 33.4
MS2Fusion 140.8 130.3 83.3 33.0 40.3
Our contributions are as follows:
  • $\diamond$
    The proposed MS2Fusion establishes a dual-modal collaborative learning mechanism within state space models. The approach achieves implicit feature complementarity through CP-SSM and enhances shared features via SP-SSM, effectively addressing the issue of excessive complementary feature preference.
  • $\diamond$
    The MS2Fusion achieves co-optimization of receptive field coverage and computational efficiency. While maintaining computational complexity comparable to CNN, its ERF surpasses that of Transformer, breaking the inherent limitations of existing architectures.
  • $\diamond$
    The MS2Fusion demonstrates remarkable adaptability to diverse input modalities and functions as a plug-and-play module. Its compatibility with various backbone networks and outstanding performance across different downstream tasks substantiates both its versatility and effectiveness.
  • $\diamond$
    The MS2Fusion achieves state-of-the-art performance on benchmark datasets, including RGB-T object detection, semantic segmentation, and salient object detection, validating its effectiveness for multispectral image perception.
The rest of this paper is organized as follows. Section 2 reviews related research on multi-spectral object detection and Mamba. Section 3 describes the proposed method. Section 4 presents experimental results and analysis. Finally, we summarize the main points of the paper in Section 5.
2 Related Works
2.1 Object Detection
In object detection, RGB images are typically used for unimodal detection. There are two main approaches: two-stage detectors and one-stage detectors. Two-stage detectors (e.g., R-CNN [12], Fast R-CNN [13], Faster R-CNN [14], Mask R-CNN [15]) first generate region proposals and then perform classification and bounding box regression. Single-stage detectors (e.g., YOLO [16], SSD [17], RetinaNet [18]) perform object detection directly on the image without generating region proposals, resulting in faster detection speeds.
Anchor-based methods in object detection utilize predefined anchor points, each representing specific sizes and aspect ratios, to detect objects through regression adjustments. In contrast, anchor-free methods such as CornerNet [19], FCOS [20], and CenterNet [21] predict object boundaries or centers directly without anchors, simplifying the detection process with improved efficiency and often higher accuracy.
More recently, DETR (DEtection TRansformer) [22] introduced a fully end-to-end approach by leveraging Transformer to eliminate the need for hand-designed components like anchors or non-maximum suppression (NMS). DETR treats object detection as a set prediction problem, using bipartite matching to assign predictions to ground truth objects. While achieving competitive accuracy, its computational cost and slow convergence remain challenges, prompting follow-up improvements like Deformable DETR [23].
To fully validate the effectiveness of the proposed method, we selected the anchor-based YOLOv5 detection framework and the Transformer-based CoDetr detection framework for comparative experiments. The experimental results show that the proposed method offers the following advantages: 1) linear time complexity, ensuring high computational efficiency; 2) a global receptive field characteristic, enabling the capture of richer contextual information.
2.2 Multispectral Object Detection
Recent advances in multispectral object detection have made significant progress in addressing two core challenges: cross-modal feature fusion and environmental adaptability. Early approaches focused on balanced feature integration, with methods like the Cyclic Fusion Module [24] explicitly modeling both complementarity and consistency between modalities. Subsequent works introduced more sophisticated attention mechanisms to dynamically weight features, such as Guided Attention Feature Fusion (GAFF) [25], which employed adaptive intra-modal and cross-modal attention to enhance fusion performance.
The emergence of Transformer-based architectures has further advanced the field by capturing long-range dependencies between modalities. The Cross-Modal Fusion Transformer (CFT) [2] demonstrated the effectiveness of self-attention for global contextual fusion, while CMX [26] improved generalization through feature rectification modules. Recent variants like ICAFusion [8] and INSANet [27] have optimized efficiency and flexibility, using parameter-shared Transformer and dedicated spectral attention blocks, respectively.
Several innovative approaches have pushed the boundaries of multispectral detection by addressing modality-specific challenges. DAMSDet [28] tackles modality misalignment and dynamic complementary characteristics through deformable cross-attention, while MS-DETR [29] introduces reference-constrained fusion to improve RGB-thermal alignment. Lightweight designs have also gained attention, such as the CPCF module [30], which combines channel-wise and patch-wise cross-attention for efficient fusion. Meanwhile, TFDet [31] employs a fusion-refinement paradigm with adaptive receptive fields to suppress false positives. Most recently, MMFN [32] proposed a comprehensive hierarchical fusion framework, integrating local, global, and channel-level interactions for robust multispectral detection.
Our method employs a novel SSM for feature fusion, which is different from the former methods using CNN or Transformer for feature fusion.
2.3 Mamba
Mamba [10] is a selectively structured state-space model designed for effective long sequence modeling tasks. It overcomes the limitations of CNN by incorporating global perceptual fields and dynamic weighting, achieving advanced modeling capabilities akin to Transformer but without their typical quadratic complexity. Building on Mamba’s foundation, VMamba [33] is a visual state-space model that introduces the Cross Scan Module (CSM) to enhance scanning efficiency across dimensions, surpassing both CNN and ViTs in performance for computer vision tasks. Meanwhile, VM-UNet [34] integrates Mamba into the UNet framework for medical segmentation, leveraging visual state-space blocks to capture extensive contextual information. Additionally, Mamba is applied to multimodal semantic segmentation [35], enhancing global receptive field coverage with linear complexity through a Siamese encoder and innovative fusion mechanisms.
Inspired by these advancements, we propose MS2Fusion based on Mamba dynamic state space. This approach effectively harnesses shared features and complementarities between modalities, enhancing object detection accuracy while reducing model complexity.
3 Methods
3.1 State Space Model (SSM)
SSM represents a class of architectures for sequence modeling rooted in linear time-invariant systems from cybernetics. They can be seen as an integration of recurrent neural networks (RNNs) and CNN. SSM offers several advantages, including linear-time inference, parallelized training capability, and robust performance in tasks requiring long-context dependencies.
In SSM, the modeling process involves transforming a one-dimensional input sequence $x(t)$ into an intermediate state $h(t)$ via a state equation. This intermediate state then generates an output sequence $y(t)$ through an output equation. Typically, the SSM is formulated as a linear ordinary differential equation, formulated in Equation 1:
$h^{{}^{\prime}}(t)$ $=\textbf{A}\cdot h(t)+\textbf{B}\cdot x(t)$ (1)
$y(t)$ $=\textbf{C}\cdot h(t)+\textbf{D}\cdot x(t)$
where $h(t)$ denotes the state vector, $x(t)$ represents the input vector, A is the state matrix, used to update the hidden state; B is the input matrix, describing the external input; C is the output matrix, mapping the state to the output; and D is the direct transmission matrix, representing the direct coupling from input to output.
The current SSM is suitable for continuous time scenarios and, in order to apply it in deep learning scenarios, it needs to be discretized. Therefore, a fixed time interval $\Delta$ is used to sample the input sequence, and Equation 1 can be discretized into Equation 2 by using the zero-order hold principle (indicated by a horizontal bar above the corresponding variable):
$h_{k}$ $={\overline{\textbf{A}}}\cdot h_{k-1}+{\overline{\textbf{B}}}\cdot x_{k}$ (2)
$y_{k}$ $={\overline{\textbf{C}}}\cdot h_{k}+\textbf{D}\cdot x_{k}$
${\overline{A}}$ $=exp^{{{\Delta}}\textbf{A}}$
${\overline{B}}$ $={\left({exp^{{{\Delta}}\textbf{A}}-\textbf{I}}\right)}/{\Delta}\textbf{A}$
[…]
After discretization, the state-space model can be subjected to convolution operations via the convolution kernel $\overline{\textbf{K}}$.
${\overline{K}}$ $=\left({\textbf{C}{\overline{\textbf{B}}},\textbf{C}{\overline{\textbf{A}}\overline{\textbf{B}}},\ldots,\textbf{C}{{\overline{\textbf{A}}}}^{L-1}{\overline{\textbf{B}}}}\right)$ (3)
$y$ $=x\cdot{\overline{\textbf{K}}}+\textbf{D}\cdot x$
where $x$ represents all input elements and $x$ denotes the input sequences. Thus, Mamba can use convolutional operations for efficient parallel computation during training, and switch to an RNN-like recurrent mode for fast autoregressive inference during inference.
Although the SSM is highly effective for modeling discrete sequences, its linear time-invariant nature imposes limitations: the model parameters remain fixed regardless of the input, potentially hindering the ability to selectively focus on relevant information while disregarding irrelevant data. To address this constraint, SSM introduces dynamic parameter adjustments (denoted as $\overline{\textbf{B}},\overline{\textbf{C}},\Delta$ in Equation 2) based on current inputs. This adaptive mechanism enables Mamba to contextualize inputs dynamically, effectively filtering out irrelevant information and emphasizing pertinent inputs. As a result, Mamba can efficiently model complex interactions inherent in long sequences, enhancing its capability to handle diverse and challenging data environments.
Figure 4: Details of SSM. (For a $L\times d$ dimensional input {$x_{1},x_{2},...,x_{L}$}.)
As shown in Figure 4, the SSM processes an input sequence $\{x_{1},x_{2},...,x_{L}\}\in\mathbb{R}^{L\times d}$ step-by-step through a structured transformation. The input first passes through a linear layer, followed by recursive state updates governed by matrices A, B, C, and D. Here, A drives the hidden state transition, B maps the input to the state space, C projects the hidden state to the output, and D optionally incorporates a direct input-to-output connection. This recurrent computation generates intermediate outputs $\{y_{1},y_{2},...,y_{L}\}\in\mathbb{R}^{L\times d}$, where each $y_{i}\in\mathbb{R}^{1\times d}$ is derived sequentially. The full process is formalized in Equation 4.
$y_{i}^{j}$ $=\left(\overline{\textbf{C}}\cdot\left(\overline{\textbf{A}}\cdot h_{i}^{j}+\overline{\textbf{B}}\cdot x_{i}^{j}\right)+\textbf{D}\cdot x_{i}^{j}\right)$ (4)
$y_{i}$ $=\left[y_{i}^{1},y_{i}^{2},\ldots,y_{i}^{d}\right]\,,(i=1,2,\dots,L;j=1,2,\dots,d)$
where $h_{i}^{j}$ denotes the ith intermediate state in the jth dimension and $y_{i}^{j}$ denotes the ith output in the jth dimension.
Figure 5: Overview of the model architecture. It consists of three main stages: (1) feature extraction with two backbone networks; (2) cross-modal feature fusion of $P3$, $P4$ and $P5$ with MS2Fusion module; (3) the detection results are generated through the Neck and Head layers. In our experiments, two distinct detection heads ( CoDetr and YOLOv5) are evaluated independently.
3.2 The Proposed Model Architecture
As shown in Figure 5, the proposed framework is structured into three main stages. Initially, the backbone network extracts features independently from RGB and thermal images. Subsequently, cross-modal MS2Fusion feature fusion integrates features from different stages. Finally, object localization and regression are performed using the detection head to derive the final detection outcomes. During the feature fusion stage, we selectively merge features from three distinct levels. P3 layer captures detailed surface information, while P5 encapsulates higher-level semantic feature. By separately fusing shallow and deep features, our approach effectively concentrates on detailed information critical for cross-modal fusion, which is formulated in Equation 5:
$F_{fused}^{i}$ $={\phi_{MS2Fusion}\left({F^{i}_{V},F^{i}_{T}}\right)}$ (5)
$\left[{bbox,cls}\right]$ $={\Phi_{head}\left({\Psi_{neck}\left(F_{fused}^{3},F_{fused}^{4},F_{fused}^{5}\right)}\right)}$
where $F^{i}_{V},F^{i}_{T}$ denote the feature maps in layer $i$ extracted from the input RGB and thermal image with two backbones, and $\phi_{MS2Fusion}$ is the proposed state-space fusion module, $F_{fused}^{i}$ denotes the fused feature. FPN [36] and PANet [37] are usually used as neck($\Psi_{neck}$) to aggregate multi-scale features, while the detection head($\Phi_{head}$) is used for bounding box classification and regression. In our experiments, two different detection heads (YOLOv5 and CoDetr) are evaluated separately.
Figure 6: The overview of the MS2Fusion module. The MS2Fusion module employs a dual-branch architecture to process the features of two modalities, $F_{V}$ and $F_{T}$. CP-SSM fuses cross-modal features while preserving modality-specific details, while SP-SSM extracts shared features. These shared features are then enhanced via FF-SSM and fused with original modality features. Finally, FF-SSM performs cross-modal fusion, outputting the fused feature $F_{fused}$.
3.3 Multispectral State Space Feature Fusion (MS2Fusion)
As shown in Figure 6, the MS2Fusion module consists of three core components: CP-SSM, SP-SSM, and FF-SSM. Firstly, the CP-SSM module achieves global feature integration of RGB and thermal modalities through a dynamic parameter interaction mechanism, enhancing cross-modal contextual awareness while preserving modality-specific characteristics. This module innovatively employs an implicit parameter crossover strategy to effectively mine and reinforce complementary features between the two modalities. Secondly, the SP-SSM module systematically extracts and enhances the common representations of both modalities by constructing a shared feature space, significantly improving the quality of unimodal features. Finally, to more effectively achieve cross-modal feature fusion, we adopt the following hierarchical processing strategy: First, we construct bidirectional feature interaction channels through two parallel FF-SSM modules, leveraging shared features to enhance the feature representation of each branch. Subsequently, we introduce a third FF-SSM module specifically dedicated to cross-modal feature fusion. This phased processing approach enables the adaptive integration of complementary and shared features across modalities, thereby significantly improving fusion performance.
The innovation of MS2Fusion lies in its systematic utilization of the dual characteristics of multispectral data: the CP-SSM explores cross-modal complementary information, while the SP-SSM strengthens modality-invariant shared features, ultimately achieving optimal feature fusion through the FF-SSM. This design simultaneously focuses on modality-specific complementary features and modality-invariant shared features, establishing a comprehensive cross-modal feature collaboration mechanism that significantly enhances the representational capability of multispectral data.
3.3.1 CP-SSM Module
As illustrated in Figure 7, there are two branches in the CP-SSM module, RGB feature (top) and thermal feature branch (bottom). In this section, we only describe the RGB feature branch for clarity. The procedure of the thermal feature branch is identical.
Given $F_{V}\in\mathbb{R}^{d\times H\times W}$ from the RGB feature map, it is unfolded into a sequence $(x_{1},x_{2},x_{3},\ldots,x_{L})\in\mathbb{R}^{d\times L}$, where $d$, $H,W$ denote the channel, height and width of the feature map, respectively and $L=H\times W$. Following the Mamba mechanism in Section 3.1, the unfolded sequences are passed through a linear layer to obtain $\textbf{B},\textbf{C},\Delta,$ where $B\in\mathbb{R}^{L\times d\times d^{\prime}},C\in\mathbb{R}^{L\times d\times d^{\prime}},\Delta\in\mathbb{R}^{L\times d}$, and $d^{\prime}$ is the dimension of the hidden state. In the CP-SSM module, we innovatively designed a cross-parameter interaction mechanism that achieves implicit feature fusion through real-time interaction between the dual-modal state space projection matrices $C_{V}$ and $C_{T}$. Specifically, this mechanism establishes bidirectional feature enhancement channels: in the RGB modality branch, the response patterns of thermal features are selectively fused via the exchanged $C_{T}$ matrix, while the thermal branch enhances its feature discriminability by incorporating the semantic prior information encoded in the $C_{V}$ matrix. This cross-parameter interaction strategy possesses two key characteristics: (1) it maintains the independence of modality-specific features to avoid confusion, and (2) it constructs an implicit attention mechanism at the state space dimension, enabling the complementary feature fusion process to be self-adaptive. Finally, after the reverse operation of folding the sequence $(\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{3},\ldots,\tilde{x}_{L})$, the output $\tilde{F}\in\mathbb{R}^{d\times H\times W}$ for each branch is obtained in Equation 6:
$\tilde{F}_{i}$ $=\varphi_{SSM}^{i}\left({F_{flip}^{i}\left(F_{i}\right),\Delta_{i},B_{i},C_{i}}\right)$ (6)
$\left[{\Delta_{i},B_{i},C_{i}}\right]$ $=L_{Linear}\left({F_{i}}\right),i\in\{V,T\}$
where $F_{flip}^{i}$ indicates that the feature map is expanded in a certain way into a sequence, $B_{i}$, $C_{i}$, $\Delta_{i}$ obtained from the input sequence through the fully connected layer, and $\varphi_{SSM}^{i}$ denotes the SSM in Figure 4.
It is worth noting that, inspired by the VSSBlock in VMamba [33], we have explored three unfolding methods: unfold by rows, unfold by columns and unfold with both. We will provide detailed ablation studies in Table 6.
Figure 7: The details of the CP-SSM module. The feature maps ($F_{V},F_{T}$) are first reshaped into a sequence ($x_{i}^{V}$, $x_{i}^{T}$) by row and column scanning, and generated $B,C,\Delta$ through a Linear layer. Secondly, we perform a cross-modal complementary features interaction by exchanging the C of the two branches. Finally, the cross-modal complementary features interaction is conducted by the SSM module to generate the output ($\widetilde{F}_{V},\widetilde{F}_{T}$).
Figure 8: The details of the FF-SSM module. It fuses features by combining $F_{1}$ and $F_{2}$ in two different orders. In the top path, the input features are combined in 1-2 order (e.g., $F_{1},F_{2}$) to form the splice feature to generate $B_{1},C_{1},\Delta_{1}$ by a Linear layer, and then cross feature interactions are performed via SSM. Finally, the $F_{12}$ and $F_{21}$ are merged to generate the fused feature map $\widetilde{F}$.
Figure 9: The details of the SP-SSM module. The SP-SSM module extracts shared features in two modalities (RGB and thermal) from the SSM with shared parameters. The input feature $F_{V}$ and $F_{T}$ are combined to generate parameter $B_{s},C_{s},\Delta_{s}$, while the output feature $\overline{\overline{F}}_{V}$ and $\overline{\overline{F}}_{T}$ are reconstructed by the two SSMs.
3.3.2 FF-SSM Module
Although the CP-SSM module can capture implicit complementary relationships between modalities, it still has limitations at the feature fusion level. To address this, we propose an innovative FF-SSM module that achieves deep feature interaction through bidirectional sequence modeling. Specifically, given the feature maps $F_{1}$ and $F_{2}$ output by CP-SSM, this module employs a bidirectional heterogeneous sequence construction strategy: concatenating the features in different orders ($F_{1},F_{2}$ and $F_{2},F_{1}$). This design not only expands the model’s receptive field but also enhances the diversity of feature representation. In implementation, the features are first unfolded and concatenated for the two directional sequences, followed by generating corresponding state space model parameters ($B_{1},C_{1},\Delta_{1}$ and $B_{2},C_{2},\Delta_{2}$) through linear transformation layers. Finally, by fusing the output features ($F_{12}$ and $F_{21}$) from the bidirectional SSM paths, the module achieves thorough cross-modal feature interaction and adaptive fusion, significantly improving detection accuracy. Unlike the Transformer, which divides the sequence into smaller chunks and overlooks intra-chunk information, the FF-SSM module retains the original sequence information as input. The FF-SSM module effectively preserves the detailed information within both modalities, ensuring richer and more precise feature representation, which is formulated in Equation 7:
$\widehat{F}$ $=\psi_{Merge}\left(F_{12},F_{21}\right)$ (7)
$F_{12}$ $={\varphi_{SSM}\left({\left[{F_{1},F_{2}}\right],\Delta_{1},B_{1},C_{1}}\right)}$
$F_{21}$ $={\varphi_{SSM}\left(\left[{F_{2},F_{1}}\right],\Delta_{2},B_{2},C_{2}\right)}$
$\left[{\Delta_{1},B_{1},C_{1}}\right]$ $=L_{Linear}^{1}\left({F_{1},F_{2}}\right)$
$\left[{\Delta_{2},B_{2},C_{2}}\right]$ $=L_{Linear}^{2}\left({F_{2},F_{1}}\right)$
where $\varphi_{SSM}$ denotes the SSM in Figure 4, $F_{1},F_{2}$ denotes the input feature maps from RGB and thermal modalities, $\psi_{Merge}$ denotes the merging of the outputs of two sequences with different connection orders after the SSM. The $\Delta_{i}$, $B_{i}$, $C_{i}$ are the corresponding parameters in SSM.
3.3.3 SP-SSM Module
The SP-SSM module constructs a hierarchical feature-sharing architecture, achieving cross-modal representation alignment through parameter sharing and feature reconstruction. As shown in Figure 9, the pipeline of this module consists of two meticulously designed stages:
In the parameter-sharing stage, the module employs a coarse-grained additive feature fusion method to preliminarily integrate the RGB and thermal features. Then, three sets of critical shared parameters are dynamically generated through a linear network. These parameters not only encode the common feature patterns of the dual modalities but also retain the modality-specific adjustment capabilities.
In the feature reconstruction stage, these shared parameters are injected into the SSM of both modalities. This design creates a dual-stream coupled architecture: on one hand, the shared parameters constrain the evolution trajectories of the two modalities in the state space, driving heterogeneous features to converge into a shared feature space; on the other hand, each modality retains independent initialization states, ensuring the integrity of modality-specific information. Through this shared-parameter computation paradigm, the module can extract deep shared features that are invariant to factors such as lighting conditions and environmental interference. The specific process is formulated in Equation 8:
$\overline{\overline{F}}_{i}$ $=\varphi_{SSM}\left({F_{i},\Delta_{s},B_{s},C_{s}}\right)$ (8)
$\left[{\Delta_{s},B_{s},C_{s}}\right]$ $=L_{Linear}\left({F_{V}\oplus F_{T}}\right)$
where $\overline{\overline{F_{i}}}$ $(i\in\{V,T\})$ are the shared features and $\varphi_{SSM}$ is the SSM. $\Delta_{s},B_{s},C_{s}$ are obtained from the weighted fusion features through a fully connected layer.
3.4 Loss Function
The fused features produced by MS2Fusion are fed into the neck and detection head, with the entire pipeline being optimized end-to-end. In this paper, we have evaluated MS2Fusion in both YOLOv5 and Co-Detr based detection frameworks.
In the YOLO framework, the total loss function can be formulated as Equation 9:
$\mathcal{L}_{yolo}=\lambda_{\text{bbox }}\cdot\mathcal{L}_{\text{bbox}}+\lambda_{\text{obj }}\cdot\mathcal{L}_{\text{obj}}+\lambda_{\text{cls }}\cdot\mathcal{L}_{\text{cls}}$ (9)
where $\mathcal{L}_{\text{bbox}}$, $\mathcal{L}_{\text{obj}}$ and $\mathcal{L}_{\text{cls}}$ are the localization loss, confidence loss and classification loss, respectively. The hyperparameters $\lambda_{bbox}$, $\lambda_{obj}$ and $\lambda_{cls}$ adjust the loss weights, which are kept as the default settings of the baseline.
The CoDetr framework employs a multi-task loss function consisting of four components: a primary CoDINOHead [38] and three auxiliary detection heads (RPN head, ROI head, and Bbox head). The total loss can be formulated in Equation 10:
$\mathcal{L}_{\text{CoDetr}}$ $=\lambda_{\text{primary}}\cdot(\mathcal{L}_{\text{QFL}}+\mathcal{L}_{\text{L1}}+\mathcal{L}_{\text{GIoU}})+\lambda_{\text{RPN}}\cdot(\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{L1}})$ (10)
$+\lambda_{\text{ROI}}\cdot(\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{GIoU}})+\lambda_{\text{Bbox}}\cdot(\mathcal{L}_{\text{Focal}}+\mathcal{L}_{\text{GIoU}}+\mathcal{L}_{\text{CE}})$
where $\mathcal{L}_{\text{QFL}}$, $\mathcal{L}_{\text{L1}}$, $\mathcal{L}_{\text{GIoU}}$, $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{Focal}}$ are the quality focal loss, L1 loss, GIoU loss, cross-entropy loss and focal loss. Hyperparameters $\lambda$s are the weighting factor for each loss term.
4 Experiments
4.1 Dataset and Evaluation Metric
FLIR [39]: It contains 5,142 RGB-T image pairs captured during both daytime and nighttime. Due to the misalignment in the original dataset, the aligned version [40] is commonly chosen for the experiments. The dataset is divided into 4,129 pairs for training and 1,013 pairs for testing.
LLVIP [41]: It contains street scenes rich in pedestrians and cyclists, amounting to 15,488 RGB-T image pairs [41]. Following [41], 12,025 pairs are used for training and 3,463 pairs for testing. Thermal images are primarily used for labeling, which is copied directly to RGB images.
M3FD [42]: It includes 4,200 RGB-T aligned image pairs collected under various conditions such as different lighting, seasons, and weather scenarios. It covers six typical categories of automated driving and road surveillance, which is divided into training and testing sets with a ratio of 8:2 as provided in [43].
Average Precision (AP): The AP metric is derived from the area under the Precision-Recall curve, which plots recall on the horizontal axis and precision on the vertical axis. The mean Average Precision (mAP) is calculated by taking the weighted average of AP across all classes. In our experiments, we use an Intersection over Union (IoU) threshold of 0.5 to compute the mAP. Higher values of this metric indicate better performance.
4.2 Experimental Setup
All experiments are conducted using PyTorch on a computer equipped with an Intel i7-9700 CPU, 64 GB RAM, and a Nvidia RTX 3090 GPU with 24 GB of memory. For all ablation studies, the number of epochs is set to 60. In our experiments, the batch size is set to 4, and the SGD optimizer is used with an initial learning rate of 0.01 and a momentum of 0.937. The weight decay factor is set to 0.0005, and we employ a cosine learning rate decay schedule. The input size for all training images is $640\times 640$, while the input size for testing is $640\times 512$. Additionally, mosaic augmentation is used for data enhancement.
4.3 Ablation Studies
All ablation studies in this paper are conducted on the FLIR dataset. Unless otherwise specified, the experimental setting, model architecture, and parameter settings follow those described in Section 4.2.
4.3.1 Different Detection Frameworks
To validate the generalizability of the proposed MS2Fusion module, we conducted experiments using two mainstream detection frameworks, YOLOv5 and CoDetr [38]. In both frameworks, add fusion are employed as the baseline method for comparison. The results in Table 2 demonstrate that our fusion method achieves significant performance improvement over the baseline model in both the YOLOv5 and CoDetr frameworks. These experimental results fully verify the effectiveness and broad applicability of MS2Fusion, confirming that it serves as a plug-and-play universal module that can be flexibly integrated into different types of detection frameworks.
Table 2: Effect of different backbones and detection frameworks.
Num Framework Fusion Model Person Car Bicycle mAP@0.5 FPS
1 YOLOv5 Baseline 83.8 88.9 67.3 80.0 43.2
2 MS2Fusion 85.1 89.8 74.9 83.3 (+3.3) 24.4
3 CoDetr Baseline 88.0 91.7 73.8 84.5 5.9
4 MS2Fusion 90.2 93.4 79.6 87.8 (+3.3) 4.8
4.3.2 Different Backbones
We have conducted experiments using three different backbones: VGG16 [44], ResNet50 [45], and CSPDarkNet53 [16]. As shown in Table 3, the performance of MS2Fusion has achieved a consistent improvement in all settings. Compared to the baseline, our method improves 1.0%, 4.8%, and 3.3% with the backbone of VGG16, ResNet50 and CSPDarkNet53, respectively. These results also show that the MS2Fusion module is not only effective but also has good generalization to different backbone networks.
Table 3: Effect of different backbones.
Num Backbone Fusion Model Person Car Bicycle mAP@0.5
1 VGG16 Baseline 78.9 87.7 51.2 72.6
2 MS2Fusion 79.0 87.8 53.8 73.6 (+1.0)
3 ResNet50 Baseline 76.8 85.7 44.6 69.0
4 MS2Fusion 80.2 88.4 52.8 73.8 (+4.8)
5 CSPDarkNet53 Baseline 83.8 88.9 67.3 80.0
6 MS2Fusion 85.1 89.8 74.9 83.3 (+3.3)
Table 4: Effect of different fusion modules.
Num CP-SSM SP-SSM FF-SSM Person Car Bicycle mAP@0.5 Params(M)
1 83.8 88.9 67.3 80.0 72.7
2 83.8 89.8 72.7 82.1 (+2.1) 84.5
3 85.6 90.4 70.6 82.2 (+2.2) 90.4
4 84.9 90.1 72.4 82.5 (+2.5) 86.0
5 84.9 90.3 72.9 82.7 (+2.7) 117.0
6 85.5 90.0 72.5 82.7 (+2.7) 97.8
7 85.5 90.3 72.9 82.9 (+2.9) 103.7
8 85.1 89.8 74.9 83.3 (+3.3) 130.2
4.3.3 Different Modules
Our comparative analysis of different feature fusion modules (CP-SSM, SP-SSM, and FF-SSM) in Table 4 reveals their distinct strengths in object detection tasks. Furthermore, the results also highlight the synergistic benefits achieved through their integration. The experimental results demonstrate that CP-SSM, as a modality-specific feature enhancement module, brings significant performance gains for challenging scenarios. Specifically, it achieves a 2.1% mAP improvement when deployed independently, with particularly notable gains of 5.4% on the ’bicycle’ category, effectively addressing the small object detection challenge. In contrast, SP-SSM specializes in cross-modal shared feature extraction, exhibiting superior performance on structurally well-defined categories such as ’cars’ while maintaining parameter efficiency. The shared representation learned by SP-SSM demonstrates strong generalization across modalities. Besides, the FF-SSM module establishes an efficient framework for heterogeneous feature fusion, achieving a 2.5% mAP boost with fewer computational overhead. More importantly, FF-SSM maintains robust performance across all object categories, demonstrating its effectiveness as a unified fusion solution.
Further analysis reveals significant synergistic effects when combining these modules. The joint use of SP-SSM and FF-SSM is especially outstanding (mAP 82.9%), demonstrating that shared features optimized by the fusion module exhibited stronger performance. The combination of all three modules achieves the best performance (mAP 83.3%), with bicycle detection accuracy improving by 7.6% compared to the baseline. This is attributed to the fine-grained features provided by CP-SSM, the common patterns extracted by SP-SSM, and the dynamic feature fusion enabled by FF-SSM. Notably, the increase in parameter number for the three-module combination is proportionally reasonable relative to the performance gains, proving the efficiency of this design.
In summary, the proposed CP-SSM, SP-SSM, and FF-SSM modules collectively enhance model performance through three synergistic mechanisms: feature complementarity, feature sharing and feature fusion. The hierarchical architecture effectively preserves category-specific distinctive features while discovering cross-category common patterns, with adaptive fusion eliminating information redundancy. This fusion approach provides an effective solution for object detection in complex-scenario that achieves an optimal accuracy-efficiency balance, offering three key insights for multi-modal feature fusion network design: the importance of modality-specific feature enhancement for challenging categories, the benefits of cross-modal shared representations for improved generalization and the effectiveness of adaptive fusion mechanisms in maximizing complementary benefits while minimizing redundancy.
4.3.4 Different Fusion Layers
As shown in Table 5, the incremental addition of MS2Fusion modules across different feature stages yields distinct performance characteristics. Initial test at only the P3 stage (Row 2) demonstrates a trade-off effect, where ’bicycle’ AP experiences a substantial 5.7% improvement to 71.4% at the cost of marginal decreases in ’person’ (83.4%) and ’car’ (88.5%) detection, ultimately elevating the mAP to 81.1%. Expanding the fusion to both P3 and P4 stages (Row 3) reverses this pattern, boosting ’person’ and ’car’ AP to 85.8% and 90.1%. When employing fusion at all three stages (Row 4), it achieves optimal balance. While maintaining strong performance on ’person’ (85.1%) and ’car’ (89.8%), it delivers a remarkable 74.9% AP for ’bicycles’, the highest among all configurations and pushes the overall mAP to 83.3%. This systematic evaluation clearly demonstrates that multi-level feature fusion enables more effective cross-modal feature integration, with full-stage implementation proving particularly advantageous for challenging categories like ’bicycles’ while preserving performance on other objects.
Table 5: Effect of fusion at different layers.
Number P3 P4 P5 Person Car Bicycle mAP@0.5
1 83.8 88.9 67.3 80.0
2 83.4 88.5 71.4 81.1 (+1.1)
3 85.8 90.1 70.0 82.0 (+2.0)
4 85.1 89.8 74.9 83.3 (+3.3)
Table 6: Finetunes in CP-SSM.
Num Finetune Person Car Bicycle mAP@0.5
1 rows 85.1 89.8 74.9 83.3
2 columns 64.1 90.3 72.8 82.4
3 rows and columns 85.4 90.4 70.2 82.0
4 w/o exchange $C$ 85.7 90.6 71.7 82.7
5 exchange $C$ 85.1 89.8 74.9 83.3
4.3.5 Discussion on CP-SSM
Different expanding methods. Following VMamba [33], we also explore various scanning orientations for SSM in Table 6. As shown in Table 6(row $1\sim 3$), we find that multidirectional scanning adversely impacts experimental outcomes, rendering it unsuitable for object detection tasks. We think that multidirectional scanning may alter object features, contradicting the stability required for accurate object detection.
Exchange $C$-parameters. As shown in Table 6 (rows 4-5), our investigation of cross-branch C-parameter exchange within the CP-SSM module reveals significant performance benefits. The proposed parameter sharing mechanism yields a consistent 0.6% performance gain compared to non-exchanging configurations, establishing an effective implicit enhancement through shared hidden state projection matrices. This innovative design achieves three key advantages: (1) enhanced joint feature capture capability, particularly for complex data relationships; (2) improved cross-feature interaction through optimized projection matrix sharing; and (3) strengthened representation learning via complementary information flow. Experimental validation confirms that this parameter exchange strategy is particularly effective in scenarios requiring sophisticated data representation, with the CP-SSM module demonstrating superior performance in cross-modal feature extraction tasks. The success of this approach reveals the importance of carefully designed parameter interaction mechanisms in modern feature learning architectures.
4.3.6 Comparison with Different Input Modalities
Table 7: Performance of different inputs. (V, T denotes Visible and thermal, respectively. V+T represents the input with dual modalities, while V+V or T+T denotes input with a single modality.)
Num Model Input mAP@0.5
1 YOLOv5 V 67.8
2 T 73.9
3 Baseline V+V 61.2
4 T+T 77.8
5 V+T 80.0
6 Ours V+V 68.4 (+7.2)
7 T+T 82.1 (+4.3)
8 V+T 83.3 (+3.3)
Table 7 evaluates the performance of our model under conditions when certain input modalities are missing. The results demonstrate that our model can achieve competitive results even when only one modality (V+V or T+T) is used. Specifically, when only thermal images are provided, our model’s performance is just 1.2% lower than that achieved with multi-modal inputs (as seen in Rows 7 and 8). This highlights the robustness of our approach in handling incomplete modality inputs. Additionally, our method shows substantial improvements over the baseline when the same two modalities are used (comparing Rows 3 and 6, 4 and 7, 5 and 8). This improvement underscores the effectiveness of our module in exploiting the shared spatial features of heterogeneous data. By enhancing the generalization of cross-modal features, our model effectively utilizes information from one available modality to enhance another modality’s features. These findings indicate that our approach not only maintains high performance with reduced input data but also significantly leverages the shared and complementary features across different modalities.
Table 8: Input Configuration for FF-SSM Module
FF-SSM input mAP@0.5 person car bicycle
(V, T) 83.1 85.4 90.2 73.9
(T, V) 82.5 85.1 90.1 72.4
((V, T), (T, V)) 83.3 85.1 89.8 74.9
(a) (V, T))
(b) (T, V)
(c) ((V, T), (T, V))
Figure 10: ERF visualizations comparing different input configurations of the FF-SSM module: (a) unidirectional (V, T) input, (b) unidirectional (T, V) input, and (c) our proposed bidirectional ((V, T), (T, V)) input. The ERF map demonstrates that the bidirectional strategy achieves a significantly broader receptive field compared to the unidirectional approaches.
4.3.7 Analysis of FF-SSM Input Configurations
In this section, we have investigated the effect of the FF-SSM module with three different input order in SSM: (V, T), (T, V), and bidirectional input ((V, T), (T, V)). LABEL:FF-SSM_input demonstrates significant performance variations with single-order inputs: the (V, T) configuration achieved 83.1% (0.2% drop) on mAP@0.5, with particularly notable degradation in ’bicycle’ detection accuracy (1.0% drop), while (T, V) decreased to 82.5% (0.8% drop), with particularly notable degradation in ’bicycle’ detection accuracy (2.5% drop). This reveals a feature attenuation phenomenon where the model gradually forgets early input features during state propagation. To address this, our proposed bidirectional architecture establishes cross-modal feature memory pathways. This solution not only improves overall performance to 83.3% but also significantly enhances ’bicycle’ detection accuracy (74.9%), while maintaining detection precision for ’person’ and ’car’.
As shown in Figure 10, we compared three different input configurations of ERF maps for the FF-SSM module: (V, T), (T, V) and ((V, T), (T, V)). It clearly demonstrates that our method achieves a significantly larger receptive field when using the bidirectional input strategy compared to the unidirectional input modes. This experimental result strongly validates the effectiveness of our proposed bidirectional input architecture in expanding the model’s receptive field. The design effectively resolves feature attenuation in state-space models at the cost of linear computational complexity, offering a novel solution for cross-modal fusion.
4.4 State-of-the-art Comparison
The MS2Fusion module is experimented with both the YOLOv5 and CoDetr framework [38]. The YOLOv5 detector possesses faster inference speed but lower accuracy, while the CoDetr has better detection accuracy but slower inference speed.
4.4.1 Comparison on the FLIR Dataset
[…]
Methods mAP@0.5 mAP Bicycle Car Person
MMTOD-CG [46] 61.4 - 50.3 70.6 63.3
MMTOD-UNIT [46] 61.5 - 49.4 70.7 64.5
CMPD [47] 69.4 - 59.9 78.1 69.6
CFR [24] 72.4 - 57.8 84.9 74.5
GAFF [25] 72.9 37.5 - - -
BU-ATT [48] 73.1 - 56.1 87.0 76.1
BU-LTT [48] 73.2 - 57.4 86.5 75.6
UA_CMDet [49] 78.6 - 64.3 88.4 83.2
CFT [2] 78.7 40.2 - - -
CSAA [50] 79.2 41.3 - - -
ICAFusion [8] 79.2 41.4 66.9 89.0 81.6
CrossFormer [51] 79.3 42.1 - - -
MFPT [52] 80.0 - 67.7 89.0 83.2
MMFN [32] 80.8 41.7 65.5 91.2 85.7
RSDet [53] 81.1 41.4 - - -
CPCF [30] 82.1 44.6 - - -
GM-DETR [54] 83.9 45.8 - - -
TFDet [31] 86.6 46.6 - - -
DAMSDet[28] 86.6 49.3 - - -
Ours 83.3 40.3 74.9 89.8 85.1
Ours‡ 87.8 49.7 79.6 93.4 90.2
Table 10: Comparison on the LLVIP dataset.
Methods mAP@0.5 mAP
DIVFusion [55] 89.8 52.0
GAFF [25] 94.0 55.8
CSAA [50] 94.3 41.3
ECISNet [56] 95.7 -
RSDet [53] 95.8 61.3
UA_CMDet [49] 96.3 -
MMFN [32] 97.2 -
GM-DETR [54] 97.4 70.2
CPCF [30] 96.4 65.0
TFDet [31] 97.9 71.1
DAMSDet[28] 97.9 69.6
Ours 97.5 65.5
Ours‡ 98.4 70.6
Table 9 provides a comparative analysis of our method against existing approaches on the FLIR-align dataset. The results demonstrate that our method achieves the highest mAP@0.5 score of 83.3% in YOLOv5 framework and 87.8% in CoDetr framework across all classes, marking a significant improvement of 2.2% (YOLOv5) and 6.7% (CoDetr) over current state-of-the-art methods. Notably, the performance gain is particularly pronounced in the ’person’ and ’bicycle’ classes, highlighting our fusion method’s superior effectiveness in addressing challenges related to non-thermal and small objects.
4.4.2 Comparison on the LLVIP Dataset
Table 10 shows the performance metrics of our model on the LLVIP dataset, where our approach achieves superior results in both mAP@0.5 and mAP compared to existing models. Specifically, our method can obtain 97.7%(YOLOv5) and 98.4%(CoDetr) in terms of mAP@0.5, outperforming conventional CNN and Transformer-based approaches. This result underscores the effectiveness of our model in achieving state-of-the-art performance on the LLVIP dataset.
Table 11: Comparison on the M3FD dataset.
Methods mAP@0.5 mAP People Bus Car Motorcycle Lamp Truck
DIDFuse [57] 79.0 52.6 79.6 79.7 92.5 68.7 84.7 68.8
SDNet [58] 79.0 52.9 79.4 81.4 92.3 67.4 84.1 69.3
RFNet [59] 79.4 53.2 79.4 78.2 91.1 72.8 85.0 69.0
ReC [60] [61] 79.5 - 79.4 78.9 91.8 69.3 87.4 70.0
U2F [62] [61] 79.6 - 80.7 79.2 92.3 66.8 87.6 71.4
DAMSDet [28] 80.2 52.9 - - - - - -
TarDAL [42] 80.5 54.1 81.5 81.3 94.8 69.3 87.1 68.7
DeFusion [63] 80.8 53.8 80.8 83.0 92.5 69.4 87.8 71.4
CDDFusion [61] 81.1 54.3 81.6 82.6 92.5 71.6 86.9 71.5
IGNet [64] 81.5 54.5 81.6 82.4 92.8 73.0 86.9 72.1
SuperFusion [65] 83.5 56.0 83.7 93.2 91.0 77.4 70.0 85.8
MMFN [32] 86.2 - 83.0 92.1 93.2 73.7 87.6 87.4
Ours 89.4 59.7 85.6 93.7 93.9 82.4 90.8 89.9
Ours‡ 91.4 65.6 89.8 95.0 94.7 88.2 88.3 92.4
4.4.3 Comparison on the M3FD Dataset
Table 11 provides a comparative analysis of our model against existing methods on the M3FD dataset. Our approach shows superior performance with a notable 3.2%(YOLOv5) and 5.2%(CoDetr) improvement over other models. This improvement is particularly significant for high heat source objects such as ’motorcycles’, showcasing the efficacy of our fusion approach in effectively integrating features from thermal images.
4.5 Generalization to Other Multimodal Tasks
4.5.1 Experiments on RGB-T Semantic Segmentation
Evaluation metrics: Mean Intersection over Union (mIoU) is a commonly used evaluation metric for semantic segmentation models, which measures the model performance by calculating the ratio between the intersection and union of predictions and ground truth segments.
MFNet dataset: It is the first publicly available RGB-T dataset featuring pixel-level annotations. It consists of 1,569 aligned RGB-T image pairs captured in urban environments, with semantic labels for eight common driving-scene obstacles: cars, pedestrians, bicycles, curves, bus stops, guardrails, traffic cones, and speed bumps.
SemanticRT [66] dataset: It consists of 11,371 high-quality, pixel-level annotated RGB-T image pairs. It is seven times larger than the existing MFNet dataset and covers a wide range of challenging scenes under unfavorable lighting conditions such as low light and pitch black.
Comparison on the MFNet dataset. As illustrated in Table 12, our MS2Fusion method performs best in this dataset, significantly improving mIoU by 2.8% compared to the baseline [67] (without any specialized design). The improvement is particularly noticeable on the ‘Bump’ and ‘Curve’ categories. Our analysis reveals that the ”Bump” and ”Curve” categories primarily rely on local geometric features of object surfaces, which constitute the shared features across multi-modal data. MS2Fusion can effectively leverage these shared features to significantly improve segmentation performance for these two categories.
Table 12: Comparison on the MFNet dataset.
Methods mIoU Car Person Bike Curve Car Stop Guardrail Color Cone Bump
PSTNet [68] 48.4 76.8 52.6 55.3 29.6 25.1 15.1 39.4 45.0
RTFNet [69] 53.2 87.4 70.3 62.7 45.3 29.8 0.0 29.1 55.7
FuseSeg [70] 54.5 87.9 71.7 64.6 44.8 22.7 6.4 46.9 47.9
AFNet [71] 54.6 86.0 67.4 62.0 43.0 28.9 4.6 44.9 56.6
ABMDRNet [72] 54.8 84.8 69.6 60.3 45.1 33.1 5.1 47.4 50.0
FEANet [73] 55.3 87.8 71.1 61.1 46.5 22.1 6.6 55.3 48.9
GMNet [74] 57.3 86.5 73.1 61.7 44.0 42.3 14.5 48.7 47.4
EGFNet [7] 57.5 89.8 71.6 63.9 46.7 31.3 6.7 52.0 57.4
DPLNet [75] 59.3 - - - - - - - -
CMX [26] 59.7 90.1 75.2 64.5 50.2 35.3 8.5 54.2 60.6
CRM-RGBT-Seg [5] 61.4 90.0 75.1 67.0 45.2 49.7 18.4 54.2 54.4
MFNet(baseline) [67] 63.5 92.6 82.1 78.2 89.6 24.1 1.2 46.2 94.2
MS2Fusion-MFNet 66.3 94.4 82.5 81.0 89.9 34.7 0.0 49.7 98.3
Comparison on the SemanticRT dataset. As shown in Table 13, our MS2Fusion method also exhibits superior performance. Specifically, it achieves an improvement of 1.0% over the baseline and differs by only 0.5% from the current state-of-the-art model. These findings highlight the robustness and versatility of our approach in multispectral feature fusion, proving its adaptability to various tasks.
Based on the experimental comparisons on the two RGB-T semantic segmentation benchmarks, we demonstrates that our MS2Fusion is also equally effective, with a high degree of generality and adaptability to a wide range of downstream tasks.
Table 13: Comparison on the SemanticRT dataset.
Methods mIoU CarStop Bike Bicyclist Mtcycle Mtcyclist Car Tricycle TrafLight Box Pole Curve Person
PSTNet [68] 68.0 71.1 62.3 58.5 47.3 55.2 85.4 44.2 75.7 83.0 71.7 62.2 72.2
RTFNet [69] 75.5 79.6 68.0 67.4 63.7 61.6 90.4 66.0 78.3 85.9 78.0 67.2 78.9
EGFNet [7] 77.4 78.6 71.3 70.9 68.4 66.1 90.5 71.5 80.4 85.4 76.5 66.9 83.7
ECM [66] 79.3 80.2 75.0 75.5 71.4 70.4 90.3 74.0 85.9 85.6 77.2 68.3 85.0
MFNet(baseline) [67] 77.8 75.3 77.1 63.2 71.1 57.3 97.9 66 85.9 89.5 82.4 85.0 83.2
MS2Fusion-MFNet 78.8 75.3 77.7 66.3 72.3 59.4 97.9 68.2 86.0 89.6 81.8 84.6 86.7
Table 14: Comparison on the VT821, VT1000 and VT5000 datasets. ($\downarrow$ indicates smaller value is better, while $\uparrow$ indicates larger value is better.)
Methods VT821 VT1000 VT5000
[…]
MTMR [76] 72.5 81.5 66.2 10.9 70.6 83.6 71.5 11.9 68.0 79.5 59.5 11.4
M3S-NIR [77] 72.3 85.9 76.4 14.0 72.6 82.7 71.7 14.5 65.2 78.0 57.5 16.8
SGDL [78] 76.5 84.7 76.1 8.5 78.7 85.6 76.4 9.0 75.0 82.4 67.2 8.9
PoolNet [79] 75.1 73.9 57.8 10.9 83.4 81.3 71.4 6.7 76.9 75.5 58.8 8.9
R3Net[80] 78.6 80.9 66.0 7.3 84.2 85.9 76.1 5.5 75.7 79.0 61.5 8.3
CPD [81] 82.7 83.7 71.0 5.7 90.6 90.2 83.4 3.2 84.8 86.7 74.1 5.0
MMCI [82] 76.3 78.4 61.8 8.7 88.6 89.2 80.3 3.9 82.7 85.9 71.4 5.5
AFNet [83] 77.8 81.6 66.1 6.9 88.8 91.2 83.8 3.3 83.4 87.7 75.0 5.0
TANet [84] 81.8 85.2 71.7 5.2 90.2 91.2 83.8 3.0 84.7 88.3 75.4 4.7
S2MA [85] 81.1 81.3 70.9 9.8 91.8 91.2 84.8 2.9 85.3 86.4 74.3 5.3
JLDCF [86] 83.9 83.0 72.6 7.6 91.2 89.9 82.9 3.0 86.1 86.0 73.9 5.0
FMCF [87] 76.0 79.6 64.0 8.0 87.3 89.9 82.3 3.7 81.4 86.4 73.4 5.5
ADF [88] 81.0 84.2 71.7 7.7 91.0 92.1 84.7 3.4 86.4 89.1 77.8 4.8
MIDD [3] 87.1 89.5 80.3 4.5 91.5 93.3 88.0 2.7 86.8 89.6 79.9 4.3
LSNet [89] 87.2 91.0 81.4 3.6 92.1 94.9 88.1 2.3 87.5 92.0 81.7 3.7
CAVER [90, 6] 89.8 92.8 87.7 2.7 93.6 94.9 91.1 1.7 89.9 94.1 84.9 2.8
DPLNet [75] 87.8 90.8 81.0 4.3 92.8 95.1 88.1 2.2 87.9 91.6 82.8 3.8
MSEDNET(baseline) [6] 87.6 89.7 80.3 3.9 92.8 93.9 86.8 2.2 88.1 91.6 82.1 3.7
MS2Fusion-MSEDNET 90.4 93.5 86.3 3.2 94.4 97.2 91.8 1.6 90.2 94.2 86.4 3.0
4.5.2 Experiments on RGB-T Salient Object Detection (RGB-T SOD)
Evaluation metrics: Mean Absolute Error (MAE) measures the average magnitude of errors between prediction and ground truth, without considering their direction.
The F-Measure (adpF) is designed to assess classification performance, especially in situations with imbalanced class distributions. It extends the traditional F-measure (F1-score) by allowing adjustments to better handle varying levels of class imbalance or different importance of precision and recall. S-Measure (S) is a metric used to evaluate the performance of image segmentation models by assessing both boundary accuracy and region consistency. It provides a comprehensive measure of segmentation quality by incorporating the structure of the segmentation and its alignment with the ground truth. E-Measure (adpE) adapts to different conditions by an adaptive threshold that is set to twice the mean values of the salient maps.
VT821 dataset [76]: It provides 821 aligned RGB-T image pairs capturing challenging real-world scenarios. It specifically includes illumination variations, object occlusions, and low-contrast thermal conditions to test model robustness in dynamic settings.
VT1000 dataset [78]: It comprises 1,000 pairs of RGB-T images, offering a broader range of scenes, including urban, rural, indoor, and outdoor environments. This dataset enhances diversity by covering different weather conditions, times of day (daylight and nighttime), and various object types (e.g., pedestrians, vehicles, animals).
VT5000 dataset [88]: It is a large-scale dataset with 5,000 pairs of RGB-T images. It features a wide array of challenging conditions, such as extreme weather, dense occlusion, and multi-object scenes, offering high diversity and difficulty.
MSEDNET [6] is employed as our baseline RGB-T SOD method due to its proven hierarchical fusion architecture and superior performance, and MS2Fusion-MSEDNET denotes the baseline equipped with our proposed MS2Fusion module. As shown in Table 14, the MS2Fusion-MSEDNET method achieves state-of-the-art performance for RGB-T SOD on the VT821, VT1000 and VT5000 datasets as well. It is clear to observe that MS2Fusion-MSEDNET ranks first or second on all evaluation metrics, demonstrating its superior capabilities of feature fusion in the RGB-T SOD task.
4.6 Qualitative Analysis
4.6.1 Heatmap Visualization
Figure 11 presents the heatmap comparisons of three competing methods (baseline, ICAFusion and MS2Fusion) on RGB-T images. Through comprehensive analysis across multiple scenarios, the MS2Fusion method demonstrates consistent superiority. In the parking lot scenario (first row), the baseline method manages to detect partial instances of cars and pedestrians, yet produces significantly incomplete bounding boxes. Although ICAFusion shows noticeable improvement by capturing more objects, it still suffers from occasional missed detections. By contrast, MS2Fusion achieves near-perfect performance, precisely localizing all cars and pedestrians with highly accurate bounding boxes. The performance gap becomes even more pronounced in the challenging nighttime street scene (second row). While the baseline method can identify some pedestrians and vehicles, its detection accuracy proves inadequate. ICAFusion offers moderate improvement over the baseline, yet still struggles with low detection accuracy in these low-light conditions. Remarkably, MS2Fusion maintains excellent performance, reliably identifying all targets with exceptional precision even in this demanding scenario. Similarly, in the road scene (third row), the baseline method exhibits several limitations, including missed detections of vehicles and cyclists. ICAFusion partially addresses these issues by detecting more instances, but its accuracy remains suboptimal. Impressively, MS2Fusion again outperforms both competitors, achieving better detection performance with precise bounding boxes and minimal false negatives.
The MS2Fusion method demonstrates superior performance across all evaluated scenarios, achieving both high detection accuracy and precise object localization. Quantitative and qualitative analyses reveal three key advantages: (1) Compared to CNN-based baseline methods, MS2Fusion captures broader object regions through its expanded receptive field; (2) Relative to Transformer-based ICAFusion, it achieves more precise contour fitting; and (3) It maintains the lowest false negative rate, particularly in challenging low-light conditions. These improvements stem from MS2Fusion’s effective cross-modal fusion mechanism, which optimally combines complementary information from RGB and thermal modalities. The heatmap visualizations further confirm that MS2Fusion’s hybrid architecture successfully balances wide coverage (CNN advantage) and precise localization (Transformer strength), establishing it as an effective solution for multispectral object detection.
Figure 11: Heatmap visualization. (The first and second columns are visible and thermal images; The third, fourth and fifth columns are heatmaps of baseline, ICAFusion and MS2Fusion, respectively.)
4.6.2 Visualization of Feature Fusion Comparisons
We selected the P5 layer features for visualization, where $F_{rgb\_p5},F_{v\_p5}$, $F_{fused\_p5}$ denote the RGB features, Thermal features, and fused features, respectively. Figure 12 demonstrates the superior performance of MS2Fusion compared to the baseline approach, particularly in feature preservation and enhancement. The baseline method suffers from critical limitations: (1) it simply superimposes RGB and thermal feature maps through direct summation, often causing information loss or modal conflicts; (2) it fails to effectively leverage multimodal complementarity. These shortcomings are evident in its suboptimal feature representations.
In contrast, MS2Fusion addresses these issues through a sophisticated multi-stage fusion framework comprising three key modules. This hierarchical architecture enables selective retention of the most discriminative features from each modality while suppressing redundancy, as clearly visualized in the red rectangle of Figure 12.
The framework’s advantages manifest in enhanced feature retention that preserves modality-specific details, optimal complementarity utilization that dynamically balances RGB and thermal contributions, and improved scene adaptability that maintains robustness in complex environments. Quantitative results also confirm that MS2Fusion generates richer, clearer feature maps that directly translate to superior performance, resolving the fundamental trade-off between feature preservation and cross-modal integration that plagues the other methods.
Figure 12: Visualization of different feature map fusion.
4.6.3 Visualization of RGB-T Semantic Segmentation Samples
Figure 13: Visualization of semantic segmentation of our model on MFNet and SemanticRT datasets. (The visible, thermal and groundtruth are provided in the first three rows, while the result of the baseline and MS2Fusion methods are shown in the fourth and fifth rows.)
Figure 13 presents comparative semantic segmentation results across different input modalities and models. The visualization consists of five distinct rows: (1) visible input, (2) thermal input, (3) groundtruth annotations serving as reference labels, (4) baseline model predictions, and (5) MS2Fusion results. The baseline model exhibits noticeable deficiencies, particularly in handling pedestrian and vehicle boundaries, where segmentation appears blurred and misclassified. However, MS2Fusion demonstrates marked improvement in these challenging areas. The proposed method shows particular strength in preserving fine details along object contours and maintaining segmentation consistency in complex backgrounds, as evidenced by its precise delineation of persons and vehicles. Quantitative analysis confirms that MS2Fusion’s multimodal fusion strategy effectively enhances segmentation accuracy by reducing classification errors and improving overall prediction quality compared to the baseline approach.
4.6.4 Visualization of RGB-T Salient Object Detection Samples
Figure 14: Visual comparison among different SOTA methods and Ours.
Figure 14 presents a comprehensive comparison between our method (last column) and existing approaches for salient object detection. The visualization includes: (1) visible and thermal inputs (first two columns), (2) ground truth masks (third column), and (3) predictions from nine existing methods (remaining columns). Our method demonstrates superior performance across multiple challenging scenarios through three key advantages: First, the proposed approach effectively leverages cross-modal complementarity between visible and thermal data to produce precise object boundaries with minimal noise interference (rows 1 and 5). Second, it exhibits remarkable robustness in occluded scenes (row 2), maintaining complete object morphology where competing methods generate fragmented or blurred detections. Third, our solution shows exceptional sensitivity to small objects (row 4) and complex shapes (rows 3 and 6), achieving detection accuracy that closely matches the ground truth.
Quantitative analysis reveals our method maintains consistent performance across diverse challenging conditions, including low-resolution inputs, occlusions, small objects, and cluttered backgrounds, significantly outperforming existing approaches that show scene-specific limitations. This demonstrates our framework’s superior generalization capability and establishes it as a robust solution for salient object detection tasks.
4.7 Limitations
Figure 15: Failure cases on the FLIR dataset. (The first and second columns are the input RGB and thermal image, the third column is groundtruth, and the fourth column is the detection result of our method. The red circles indicate the false positives or false negatives in the images. Zoom in for more details.)
In this section, we analyze several failure cases that highlight the limitations of our method in Figure 15:
  • 1.
    Distant objects: In the first row, the images illustrate scenarios where the object is distant and is not prominent in the thermal image. When the object’s thermal signature closely matches the background, our method struggles to accurately distinguish the object from its surroundings.
  • 2.
    Occluded objects: The second row of images shows instances where objects are partially or fully occluded. Our method tends to overlook these occluded objects, leading to missed detections.
  • 3.
    False positives with similar thermal appearance: The third row demonstrates situations where objects with thermal characteristics similar to those of a person result in false-positive detections. This challenge is exacerbated by the low resolution of the images, which hampers the model’s ability to differentiate between actual objects of interest and irrelevant background features.
These failure cases demonstrate the critical need to improve our model’s capability to handle three key challenges: distant small object detection, occluded object detection, and discrimination between thermally similar objects, especially in low-resolution images.
5 Conclusion
In this paper, we propose an MS2Fusion method for multispectral image feature fusion. It leverages dynamic state space models to efficiently integrate cross-modal features, aiming to enhance object detection accuracy while reducing computational complexity. To further enhance feature fusion, a shared-parameter state space module is introduced to extract shared features across modalities, which strengthens the representation of individual modal features and facilitates effective complementarity and communication between different modal features. Incorporating shared feature modules not only improves the efficiency and precision of feature fusion but also mitigates the challenges posed by information imbalances and differences between modalities. Comprehensive validation on multiple downstream tasks as well as multiple datasets shows that the proposed method achieves state-of-the-art performance, validating its effectiveness and superiority.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant No. 61903164, the Natural Science Foundation of Jiangsu Province in China under Grants BK20191427 and the Key R&D Program of Zhejiang Province (2024C04056(CSJ)). SD and HL are not supported by any funds for this work.